OpenStudy (anonymous):

(D-1)x + Dy = 0 (D+1)x + (2D+2)y = 0 How can i solve these equ. in MATLAB ?

OpenStudy (anonymous):

To solve these equations in Matlab, you'll need to write them a bit differently. Try to write them as a matrix-vector multiplication. Someting like \[A \cdot \left[ \begin{matrix}x \\ y\end{matrix} \right] = \left[ \begin{matrix} 0 \\ 0 \end{matrix} \right]\] Equations in that form can be solved by Matlab easily.